OpenStudy (anonymous):

The base of a 15-foot ladder is being pulled away from a wall at a rate of 11 feet per second. How fast is the top of the ladder sliding down the wall at the instant when the base of the ladder is 9 feet from the wall? ft/s=?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[x^2+y^2=15^2\] \[2xx'+2yy'=0\]you know \(x'=11,x=9,y=12\) solve for \(y'\)